OpenStudy (anonymous):

The second of three numbers is 4 times the first number. The third number is 12 less than the first. Find the three numbers given that their sum is 42.

OpenStudy (anonymous):

just forming a matrix

OpenStudy (anonymous):

let the first number be x second be y, third be z

OpenStudy (anonymous):

y= 4x z=x-12 x+y+z=42

OpenStudy (anonymous):

solve etc

OpenStudy (anonymous):

sub the first and second eqn into the third

OpenStudy (anonymous):

6x -12 = 42

OpenStudy (anonymous):

etc , its easy from now

OpenStudy (anonymous):

6x=54 x= 9

OpenStudy (anonymous):

y= 36

OpenStudy (anonymous):

z=-3

OpenStudy (anonymous):

let's call the numbers x, y, and z. We can turn the statements into equations: 1. y = 4x 2. z = x - 12 3. x + y + z = 42 Now we can put "4x" in place of "y" and "x - 12" in place of "z" in equation 3, because equations 1 and 2 tell us they are equivalent. Then we solve for x: x + 4x + x - 12 = 42 6x - 12 = 42 6x = 54 x = 9 Plug x into the first 2 equations to solve for y and d: y = 4(9) y = 36 z = 9 - 12 z = -3

OpenStudy (anonymous):

or "what elecengineer said." BTW, I'm also an electrical engineer. :-)